# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v2

@@ Abstract
Large language models (LLMs) have shown amazing performance on tasks that require planning and reasoning. Motivated by this, we investigate the internal mechanisms that underpin a network’s ability to perform complex logical reasoning. We first construct a synthetic propositional logic problem that serves as a concrete test-bed for network training and evaluation. Crucially, this problem demands nontrivial planning to solve, but we can train a small transformer to achieve perfect accuracy. Building on our set-up, we then pursue an understanding of precisely how a three-layer transformer, trained from scratch, solves this problem. We are able to identify certain “planning” and “reasoning” circuits in the network that necessitate cooperation between the attention blocks to implement the desired logic. To expand our findings, we then study a larger model, Mistral 7B. Using activation patching, we characterize internal components that are critical in solving our logic problem. Overall, our work systemically uncovers novel aspects of small and large transformers, and continues the study of how they plan and reason.
@@ 1 Introduction
While the study of how transformers reason in general remains a daunting task, in this work, we aim to improve our mechanistic understanding of how a Transformer reason through simple propositional logic problems. For concreteness’ sake, consider the following problem:
@@
The reasoning problem, while simple-looking on the surface, requires the model to perform several actions that are essential to more complex reasoning problems, all without chain of thought (CoT). Before writing down any token, the model has to first discern the rule which is being queried: in this case, it is “A or B implies C”. Then, it needs to rely on the premise variables A and B to the locate the relevant facts, and find “A is true” and “B is false”. Finally, it needs to decide that “A is true” is the correct one to invoke in its answer due to the nature of disjunction. It follows that, to write down the first token “A”, the model already has to form a “mental map” of the variable relations, value assignments and query! Therefore, we believe that this is close to the minimal problem to examine how a model internalizes and plans for solving a nontrivial mathematical reasoning problem where apparent ambiguities in the problem specification cannot be resolved trivially.
To understand the internal mechanisms of how a transformer solves problems resembling the minimal form above, we perform two flavors of experiments. The first is on shallow transformers trained purely on the synthetic propositional logic problems. This enables a fine-grained analysis in a controlled setting. The other set of experiments are on a pre-trained LLM (Mistral-7B), where we primarily rely on activation patching to uncover necessary circuits for solving the reasoning problem, including specialized roles of certain components.
At a high level, we make the following discoveries based on our two fronts of analysis:
@@
    We discover that small transformers, trained purely on the synthetic problem, utilize certain “routing embeddings” to significantly alter the information flow of the deeper layers when solving different sub-categories of the reasoning problem. We also characterize the different reasoning pathways: we find that problems querying for reasoning chains involving logical operators typically require greater involvement of all the layers in the model.
@@
Additionally, we define the scope of our analysis as follows. First, in the shallow transformer experiments, we focus on the variant which only has self-attention layers in addition to layer normalization, positional encoding, embedding and softmax parameters. While we could have also included MLP layers, we choose not to because the no-MLP models already achieve 100% accuracy on the problem, and adding MLPs would unnecessarily complicate the analysis. As a second way to focus the scope of paper, in the Mistral-7B experiments, we do not seek to uncover every model component that participates in solving the reasoning problem. We focus more on finding and analyzing the components that are necessary to the model’s reasoning circuit, and necessary towards implementing the reasoning pathway as described before. By doing so, we can fully justify the necessity of these key components, without guessing about the roles of less-impactful sub-circuits.
@@ 1.1 Related works
Mechanistic interpretability. Our work falls in the area of mechanistic interpretability, which aims to understand the mechanisms that enable capabilities of the LLM; such studies involve uncovering certain “circuits” in the network (Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2022; Vig et al., 2020; Feng & Steinhardt, 2024; Wu et al., 2023; Wang et al., 2023; Hanna et al., 2024; Merullo et al., 2024; McGrath et al., 2023; Singh et al., 2024; Feng et al., 2024). While the definition of a “circuit” varies across different works, in this paper, our definition is similar to the one in Wang et al. (2023): it is a collection of model components (attention heads, neurons, etc.) with the “edges” in the circuit indicating the information flow between the components in the forward pass; the “excitation” of the circuit is the input tokens.
Evaluation of reasoning abilities of LLMs. Our work is also related to the line of work which focus on empirically evaluating the reasoning abilities of LLMs across different types of tasks (Xue et al., 2024; Chen et al., 2024; Patel et al., 2024; Morishita et al., 2023; Seals & Shalin, 2024; Zhang et al., 2023; Saparov & He, 2023; Saparov et al., 2024; Luo et al., 2024; Han et al., 2024; Tafjord et al., 2021; Hendrycks et al., 2021; Dziri et al., 2024; Yang et al., 2024). While these studies primarily benchmark their performance on sophisticated tasks, our work focuses on understanding “how” transformers reason on logic problems accessible to fine-grained analysis.
Analysis of how LLMs reason. There are far fewer studies that focus on providing fine-grained analysis of how LLMs reason. To the best of our knowledge, only a handful of works, such as Brinkmann et al. (2024); Xue et al. (2024); Zečević et al. (2023); Ye et al. (2024), share similar goals of understanding how transformers perform multi-step reasoning through detailed empirical or theoretical analysis. However, none studies the [Variable relationships]+[Variable value assignment]+[Query] type problem in conjunction with analysis on both small transformers trained purely on the synthetic problem, and large language models trained on a large corpus of internet data.
@@ 2 Data model: a propositional logic problem
In this section, we describe the synthetic propositional logic problem that shall be the data model of this paper. Our problem follows an implicit causal structure, as illustrated in Figure 1. The structure consists of two distinct chains: One containing a logical operator at the end of the chain, and the other forming a purely linear chain.
Figure 1: Synthetic data model. The causal structure has two chains: one with a logical operator (LogOp) at the end and the other being purely a linear causal chain. This example is the length-3 case and all symbols are explained in Section 2.
@@
Importance of the first answer token. Correctly writing down the first answer token is central to the accuracy of the proof, because as discussed in Section 1, it requires the model to mentally process every part of the context properly due to the minimal-proof requirement of the solution.
3 Mechanisms of planning and reasoning: a case study of the length-3 problem
@@
We pay particular attention to the first question, because as noted in Section 2, the first answer token reveals the most about how the transformer mentally processes all the context information without any access to chain of thought (CoT). We delay the less interesting answer of question 2 to the Appendix due to space limitations.
@@ 3.1 Learner: a decoder-only attention-only transformer
Architecture choice for mechanistic analysis. We select a 3-layer 3-head transformer to initiate our analysis since it is the smallest transformer that can achieve 100% accuracy; we also show the accuracies of several candidate model sizes in Figure 7 in Appendix B for more evidence. Note that a model’s answer on a problem is considered accurate only if every token in its answer matches that of the correct answer.
@@ 3.2.1 Linear or LogOp chain: routing signal at the QUERY position
Evidence 1b: distinct layer-3 attention behavior w.r.t. chain type. When the linear chain is queried, the layer-3 attention heads predominantly focus on the QUERY position, with over 90% of their attention weights on the QUERY position on average (based on 1k test samples). In contrast, when the LogOp chain is queried, less than 5% of layer-3 attention is on the QUERY on average. Instead, attention shifts to the Rules and Facts sections of the context.
@@
3.2.2 Answer for the linear chain
At this point, it is clear to us that, when QUERY is for the linear chain, the third layer mainly serves a simple “message passing” role: it passes the information in the layer-2 residual stream at the QUERY position to the ANSWER position. One natural question arises: does the input to the third layer truly contain the information to determine the first token of the answer, namely the starting node of the linear chain? The answer is yes.
Evidence 2: linearly-decodable linear-chain answer at layer 2. We train an affine classifier with the same input as the third attention block at the QUERY position, with the target being the start of the linear chain; the training samples only query for the linear chain, and we generate 5k of them. We obtain a test accuracy above 97% for this classifier (on 5k test samples), confirming that layer 2 already has the answer at the QUERY position.
3.2.3 Answer for the LogOp chain
@@
[…]
@@
4 The reasoning circuit in Mistral-7B
We now turn to examine how a pretrained LLM, namely Mistral-7B-v0.1, solves this reasoning problem. We choose this LLM as it is amongst the smallest accessible model which achieves above 70% accuracy on (a minimal version of) our problem. Our primary focus here is the same as in the previous section: how does the model infer the first answer token without any CoT? We are interested in this question as the first answer token requires to model to process all the information in the context properly without access to any CoT.
We describe the main properties of the reasoning circuit inside the model for this prediction task in Figure 3. At a high level, there are several intriguing properties of the reasoning circuit of the LLM: [^1] We use $(\ell,h)$ to denote an attention head. When referencing multiple heads in the same layer, we write $(\ell,h_{1};h_{2};...;h_{n})$ for brevity.
@@ 4 The reasoning circuit in Mistral-7B
    There is a sparse set of attention heads that are found to be central to the reasoning circuit:
    • (Queried-rule locating head) Attention heads (9,25;26), (12,9), (14,24;26) locate the queried rule using the QUERY token, and stores this information at the QUERY position.
    • (Queried-rule mover head) Attention heads (13,11), (15,8) move QUERY and the queried-rule information from the QUERY position to the “:” position.
    • (Fact processing heads) Attention heads (16,12;14) locate the relevant facts, and move information to the “:” position.
    • (Decision head) Attention heads (19,8;9;16), (17,25), relying on the aggregated information, makes a decision on which token to output.
@@ 4.1 Minimal problem description
In our Mistral-7B experiments, the input samples have the following properties:
@@
    We give the model 6 (randomly chosen) in-context examples before asking for the answer.
@@
The design decision in the first point is to ensure fairness to the LLM which was not trained on our specific logic problem. As for the last two point, we restrict the problem in this fashion mainly to ensure that the first answer token is unique, which improves the tractability of the analysis. Note that these restrictions do not take away the core challenge of this problem: the LLM still needs to process all the context information without CoT to determine the correct first token.
@@ 4.2 Causal mediation analysis
We provide evidence in this part of the paper primarily relying on a popular technique in mechanistic interpretability: causal mediation analysis. Our methodology is roughly as follows:
@@
    $\Delta_{orig,n}=\text{logit}({\bm{X}}_{orig,n})[y_{orig,n}]-\text{logit}({\bm{X}}_{orig,n})[y_{alt,n}].$
    For a high-accuracy model, $\Delta_{orig,n}$ should be large for most $n$’s, since it must be able to clearly tell that on an ${\bm{X}}_{orig,n}$, it is the LogOp chain which is being queried, not the linear chain.
@@
    Now average the $\Delta_{orig\to alt,n;(\ell,h,t)}$’s over $n$ for every $\ell,h$ and $t$ (recall that $n$ is the sample index).
@@
    This procedure helps us identify components that are significant in processing and passing the QUERY information for inference. Intuitively, an activation that result in a positive and large $\Delta_{orig\to alt,n;(\ell,t)}$ play a significant role in this subtask, because this activation helps “altering” the model’s “belief” from “QUERY is for the LogOp chain” to “QUERY is for the linear chain”.
@@
    Remark: due to the symmetry of this running example, it is perfectly sensible to perform $alt\to orig$ interventions too, by mirroring the above procedures.

### v3

@@ Abstract
Large language models (LLMs) have shown amazing performance on tasks that require planning and reasoning. Motivated by this, we investigate the internal mechanisms that underpin a network’s ability to perform complex logical reasoning. We first construct a synthetic propositional logic problem that serves as a concrete test-bed for network training and evaluation. Crucially, this problem demands nontrivial planning to solve. We perform our study on two fronts. First, we pursue an understanding of precisely how a three-layer transformer, trained from scratch and attains perfect test accuracy, solves this problem. We are able to identify certain “planning” and “reasoning” mechanisms in the network that necessitate cooperation between the attention blocks to implement the desired logic. Second, we study how pretrained LLMs, namely Mistral-7B and Gemma-2-9B, solve this problem. We characterize their reasoning circuits through causal intervention experiments, providing necessity and sufficiency evidence for the circuits. We find evidence suggesting that the two models’ latent reasoning strategies are surprisingly similar, and human-like. Overall, our work systemically uncovers novel aspects of small and large transformers, and continues the study of how they plan and reason.
@@ 1 Introduction
While the study of how transformers reason in general remains a daunting task, in this work, we aim to improve our mechanistic understanding of how a transformer reasons through simple propositional logic problems. For concreteness’ sake, consider the following problem:
@@
The reasoning problem, while simple-looking on the surface, requires the model to perform several actions that are essential to more complex reasoning problems, all without chain of thought (CoT). Correctly writing down the first answer token alone takes multiple steps. It must first resolve the ambiguity of which rule is being queried relying on the question: in this case, it is “A or B implies C”. Then, it needs to rely on the premise variables A and B to resolve the ambiguity of which facts are relevant, and identify “A is true” and “B is false”. Finally, it needs to decide that “A is true” is the correct one to invoke in its answer due to the nature of disjunction.
[…]
Contributions. At a high level, we make the following discoveries based on our two fronts of analysis:
@@
    3) We discover that small transformers, trained purely on the synthetic problem, utilize certain “routing embeddings” to significantly alter the information flow of the deeper layers when solving different sub-categories of the reasoning problem. We also characterize the different reasoning pathways: we find that problems querying for reasoning chains involving logical operators typically require greater involvement of all the layers in the model.
@@
In particular, our analysis of Mistral-7B and Gemma-2-9B in §4 and §C.7 is, to our knowledge, the first to characterize the circuits employed by LLMs in the wild for solving a nontrivial in-context logical reasoning problem (involving distracting clauses) that requires correctly executing a multi-step reasoning chain without chain of thought. Furthermore, we make concrete progress towards general techniques of verifying the necessity and sufficiency of a pretrained LLM’s reasoning circuit.
Additionally, we define the scope of our analysis as follows. First, in the shallow transformer experiments, we focus on the variant which only has self-attention layers in addition to layer normalization, positional encoding, embedding and softmax parameters. While we could have also included MLP layers, we choose not to because the no-MLP models already achieve 100% test accuracy on the problem, and adding MLPs would unnecessarily complicate the analysis. As a second way to focus the scope of paper, in the LLM experiments, we do not seek to uncover every model component that participates in solving the reasoning problems. Instead, we focus on those which have the strongest causal influence on the model’s reasoning actions. By doing so, we can fully justify the necessity of these key components for the model’s reasoning actions, without diluting the focus with less-impactful sub-circuits (for handling low-level text processing, for instance).
@@ 1.1 Related works
Mechanistic interpretability. Our work falls in the area of mechanistic interpretability, which aims to understand the hidden mechanisms that enable language-modeling capabilities of the LLMs (Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2022; Vig et al., 2020; Feng & Steinhardt, 2024; Wu et al., 2023; Wang et al., 2023; Hanna et al., 2024; Merullo et al., 2024a; McGrath et al., 2023; Singh et al., 2024; Geiger et al., 2024; Feng et al., 2024; Marks et al., 2024; Engels et al., 2024; Merullo et al., 2024b). In particular, our work belongs to the sub-area of circuit-based mechanistic interpretability (Elhage et al., 2021; Wang et al., 2023). While the definition of a “circuit” varies across different works, in this paper, our definition is similar to the one in Wang et al. (2023): it is a collection of model components (attention heads, neurons, etc.) with the “edges” in the circuit indicating the information flow between the components in the forward pass; the “excitation” of the circuit is the input tokens.
Evaluation of reasoning abilities of LLMs. Our work is also related to the line of work which focus on empirically evaluating the reasoning abilities of LLMs across different types of tasks (Xue et al., 2024; Tafjord et al., 2021; Hendrycks et al., 2021; Chen et al., 2024; Patel et al., 2024; Berglund et al., 2024; Morishita et al., 2023; Fu et al., 2024; Seals & Shalin, 2024; Zhang et al., 2023; Saparov & He, 2023; Saparov et al., 2024; Luo et al., 2024; Arora et al., 2024; Han et al., 2024; Dziri et al., 2024; Yang et al., 2024). While these studies primarily benchmark their performance on sophisticated tasks, our work focuses on understanding “how” transformers reason on logic problems accessible to fine-grained analysis.
Analysis of how LLMs reason. There are far fewer studies that focus on providing fine-grained analysis of how LLMs reason. To the best of our knowledge, only a handful of works, such as Wang et al. (2023); Zečević et al. (2023); Ye et al. (2024), share similar goals of understanding how transformers perform multi-step reasoning through detailed empirical or theoretical analysis. However, none studies the [Variable relationships]+[Variable value assignment]+[Query] type problem in conjunction with analysis on both small transformers trained purely on the synthetic problem, and large language models trained on a large corpus of internet data. Moreover, there is little study on the existence and properties of circuits employed by LLMs (in the wild) for performing latent in-context multi-hop reasoning.
@@ 2 Data model: a propositional logic problem
Figure 1: Synthetic data model. The causal structure has two chains: one with a logical operator (LogOp) at the end and the other being purely a linear causal chain. They do not share any proposition variable. This example is the length-3 case and all symbols are explained in §2.
In this section, we describe the synthetic propositional logic problem that shall be the data model of this paper. Our problem follows an implicit causal structure, as illustrated in Figure 1. The structure consists of two distinct chains: exactly one containing a logical operator at the end of the chain, and one forming a purely linear chain. The two chains do not share any proposition variable.
@@
Importance of the first answer token. As discussed in §1, correctly writing down the first answer token is central to the accuracy of the proof. First, it requires the model to process every part of the context properly without CoT due to the minimal-proof requirement of the solution. Moreover, it is the answer token position which demands the greatest number of latent reasoning steps in the whole answer, making it the most challenging token for the model to resolve. Therefore, this token is the most interesting place to focus our study on, as we are primarily interested in how the model internalizes and plans for the reasoning problems in this work.
3 The reasoning circuit in a small transformer
@@
We pay particular attention to the first question, because as noted in §2, the first answer token reveals the most about how the transformer mentally processes all the context information without any access to chain of thought (CoT). We delay the less interesting answer of question 2 to the Appendix due to space limitations.
@@ 3.1 Learner: a decoder-only attention-only transformer
Architecture choice for mechanistic analysis. We select a 3-layer 3-head transformer to initiate our analysis since it is the smallest transformer that can achieve 100% test accuracy; we also show the accuracies of several candidate model sizes in Figure 6 in Appendix B for more evidence. Note that a model’s answer on a problem is considered accurate only if every token in its answer matches that of the correct answer. Please refer to Appendix B.2 for an illustration of the model components.
@@ 3.2 Mechanism analysis
We illustrate the overall reasoning strategy and core evidence for it in Figure 8 in Appendix B.3.
@@ 3.2.1 Linear or LogOp chain: routing signal at the QUERY position
Evidence 1b: distinct layer-3 attention behavior w.r.t. chain type. When the linear chain is queried, the layer-3 attention heads predominantly focus on the QUERY position, with over 90% of their attention weights on the QUERY position on average (based on 1k test samples). In contrast, when the LogOp chain is queried, less than 5% of layer-3 attention is on the QUERY on average. Instead, attention shifts to the Rules and Facts sections of the context, as shown in Figure 9 in Appendix B.4.
@@
3.2.2 Answer for the linear and LogOp chain
Linear chain: answer at layer-2 residual stream at QUERY position. At this point, it is clear to us that, when QUERY is for the linear chain, the third layer mainly serves a simple “message passing” role: it passes the information in the layer-2 residual stream at the QUERY position to the ANSWER position. One natural question arises: does the input to the third layer truly contain the information to determine the first token of the answer, namely the starting node of the linear chain? The answer is yes.
Evidence 2: linearly-decodable linear-chain answer at layer 2. We train an affine classifier with the same input as the third attention block at the QUERY position, with the target being the start of the linear chain; the training samples only query for the linear chain, and we generate 5k of them. We obtain a test accuracy above 97% for this classifier (on 5k test samples), confirming that layer 2 already has the answer linearly encoded at the QUERY position. To add further contrasting evidence, we train another linear classifier with exactly the same task as before, except it needs to predict the correct start of the LogOp chain. We find that the classifier achieves a low test accuracy of approximately 27%, and exhibits severe overfitting with the training accuracy around 94%.
@@ 3.2.2 Answer for the linear and LogOp chain
We delay the full set of evidence for the above two claims to Appendix B.4 due to space limits. Our argument mainly relies on linear probing and causal interventions at different layers and token positions in the model.
4 The reasoning circuit in LLMs
We now turn to examine how pretrained LLMs, namely Mistral-7B-v0.1 and Gemma-2-9B, solves this reasoning problem. We choose these LLMs as they are amongst the smallest accessible models which achieves above 70% accuracy on (a minimal version of) our problem. Our primary focus here is the same as in the previous section: how does the model infer the first answer token without any CoT? We are interested in this question as the first answer token requires to model to process all the information in the context properly without access to any CoT, and demands the greatest number of latent reasoning steps out of all answer tokens.
@@ 4 The reasoning circuit in LLMs
In the main text, we primarily focus on presenting the reasoning circuit discovery and verification results of Mistral-7B, and delay the results for Gemma-2-9B to Appendix C.7. We illustrate the reasoning circuit inside Mistral-7B for this prediction task in Figure 3. At a high level, there are two intriguing properties of the reasoning circuit of the LLM: [^1] We use $(\ell,h)$ to denote an attention head. When referencing multiple heads in the same layer, we write $(\ell,h_{1};h_{2};...;h_{n})$ for brevity.
@@
    There is a sparse set of attention heads that are found to be central to the reasoning circuit: the queried-rule locating heads, queried-rule mover heads, fact-processing heads, and decision heads. We discuss circuit discovery in §4.3.1, and circuit verification in §4.3.2. We summarize the high-level conclusions of our more involved experiments in §4.3.3, and refer the reader to the respective sections in the Appendix from there.
@@ 4.1 Minimal problem description
In our LLM experiments, the input samples have the following properties:
@@
    Before asking for the answer, we give Mistral-7B 6 randomly chosen in-context examples, and 4 for Gemma-2-9B.
@@
The design decision in the first point is to ensure fairness to the LLM which was not trained on our specific logic problem. As for the last two points, we restrict the problem in this fashion mainly to ensure that the first answer token is unique, which improves the tractability of the analysis. Note that these restrictions do not take away the core challenge of this problem: the LLMs still need to process all the context information without CoT to determine the correct first answer token, and out of all answer token positions, the first token still requires the greatest number of latent reasoning steps to resolve. We discuss concrete examples and finer details of testing the models on this problem in Appendix C.
@@ 4.2 Causal mediation analysis
We provide evidence in this part of the paper primarily relying on a popular technique in mechanistic interpretability: causal mediation analysis. Our methodology is as follows:
@@
    $\Delta_{orig,n}=\text{logit}({\bm{X}}_{orig,n})[y_{orig,n}]-\text{logit}({\bm{X}}_{orig,n})[y_{alt,n}],$ (1)
    $\Delta_{alt,n}=\text{logit}({\bm{X}}_{alt,n})[y_{alt,n}]-\text{logit}({\bm{X}}_{alt,n})[y_{orig,n}].$
    For a high-accuracy model, $\Delta_{orig,n}$ and $\Delta_{alt,n}$ should be large and positive for most $n$’s, since most of the time it must be able to clearly tell that on an ${\bm{X}}_{orig,n}$, it is the LogOp chain which is being queried, not the linear chain (similarly on an ${\bm{X}}_{alt,n}$).
@@
    Average the $\Delta_{orig\to alt,n;(\ell,h,t)}$’s over $n$ for every $\ell,h$ and $t$ (recall that $n$ is the sample index).
@@
    This procedure helps us identify components that are significant in processing the QUERY information for inference. Intuitively, an activation that result in a positive and large $\Delta_{orig\to alt,n;(\ell,t)}$ (with an ideal upper limit being $\Delta_{alt,n}$) play a significant role in this subtask, because this activation helps “altering” the model’s “belief” from “QUERY is for the LogOp chain” to “QUERY is for the linear chain”. We denote the components identified as a set $\mathcal{C}$.
@@
    Remark: due to the symmetry of this running example, it is perfectly sensible to perform $alt\to orig$ interventions too, by mirroring the above procedures; we indeed adopt this mirrored procedure too in our experiments for efficient use of model computations.
QUERY-based activation patching is important to reasoning circuit discovery and verification. To justify this claim, there are two points to emphasize first. (1) To solve the reasoning problem, the QUERY is critical to initiating the reasoning chain: without it, the rules and facts are completely useless; with it, the reasoner can then proceed to identify the relevant rules and facts to predict the answer. (2) The prompt pairs differ only by the QUERY token: there is complete information for the model to provide answer for both chains given the QUERY.
If performing the aforementioned interventions on a model component leads to a large intervened logit difference (i.e. it alters the model’s “belief”), then this component must be integral to the reasoning circuit, because the component is now identified to be QUERY-sensitive and has causal influence on (parts of) the model’s reasoning actions. Furthermore, if on (most of) the prompt pairs, by patching all the circuit components in $\mathcal{C}$ simultaneously while freezing the rest causes the intervened logit difference to approach the “maximal” altered logit difference [^3] In this case, we simply consider the “maximal” altered logit difference to be the logit difference of the model run on the altered prompts without interventions, i.e. the $\Delta_{alt,n}$’s., then we obtain evidence suggesting sufficiency of $\mathcal{C}$ for explaining the LLM’s (QUERY-sensitive) reasoning actions.
For greater clarity, we provide illustrations of the intervention experiments in Appendix C.2.